# Meridional Circulation Streamlined

### v1

@@ 4 Applications
Let us consider a radiative zone that is initially under solid body rotation, $\Omega$. At time $t=0$, a zonal wind, $u_{F}$, at the outer boundary ($z=0$) starts torquing this stably stratified layer. A fast, purely dynamical, adiabatic/geostrophic adjustment sets up meridional circulation in a layer, the depth of which depends on the ratio of rotation and stratification $\Omega/N$, only. The circulation advects the background stratification to establish a meridional temperature gradient. Section 3 of Spiegel and Zahn (1992) gives a detailed description of the process.
Our solution aims to model the subsequent much slower evolution. The meridional temperature gradient developed in the adjusted layer (see Figure 1d panel 2) also impacts the vertical temperature profiles. Following a rapid thermal adjustment occurring on a short thermal timescale (Spiegel and Zahn, 1992), radiative diffusion propagates the horizontal temperature inhomogeneities further down; Figure 1d panel 3. As a result, the zonal wind and the meridional circulation spread deeper into the radiative envelope (Section 4 Spiegel and Zahn (1992)). The Rayleigh frictional drag (or viscosity) opposes this by acting on the vertical shear of the zonal wind. In steady-state, the meridional circulation penetrates to a depth until these two effects balance each other out. This slower process depends on the Prandlt number in addition to rotation and stratification, as discussed in §2 for the steady-state meridional streamfunction.
Figure 2: Analytical solution for steady-state zonal wind varying with the vertical coordinate indicating the depth of penetration of the meridional circulation. Different values of $k$ (black) include the effects of rotation $f$, stratification $N$, meridional wavenumber $l$, scale height $H$ and the radiative relaxation timescale $\tau_{R}$ according to $k=1/(L_{D}^{2}l^{2}\tau_{R})$. The effect of the viscous drag timescale $\tau_{F}$, on the depth of penetration is shown by red lines. The black dashed line overlapping the red solid line, on the left, depicts the depth of meridional circulation in the radiative zone of the Sun to approximately extend to 2.5 scale heights below the top boundary ($z=0$), the tachocline.
@@ 4.1 Steady-state zonal velocity
Here, $k$ accounts for the radiative timescale, rotation, and stratification via $\tau_{R}$, $f$, and $N$, respectively. Figure 2 shows vertically varying steady-state zonal wind profiles for different values of $k$ (black, constant $\tau_{F}=10^{13}\hskip 2.84526ptyrs=3\times 10^{20}s$) and $\tau_{F}$ (red, constant $k=5\times 10^{-19}\hskip 2.84526ptcm^{2}/s$) using expression (16). These are values adopted as representative of the Sun (Menou et al., 2004; Menou and Le Mer, 2006), the zonal velocity profile for which is labeled in the figure (black dashed, red solid lines). The radiative-convective interface is at $z=0$, and the bottom of the domain is at $z=\ ln(10^{3})=7$, considering the passive radiative zone to extend through pressures varying three orders in magnitude. The forcing zonal velocity, $u_{F}$, is $10\hskip 2.84526ptcm/s$ without loss of generality. The bottom boundary shears freely. Thus, there is no hard constraint on the velocity, allowing it to assume any value.
@@
From the same figure, the zonal wind is seen to penetrate all the way through the radiative zone of the Sun at steady state. This is in reasonable agreement with the steady-state estimate made using equation (5.19) of Spiegel and Zahn (1992) by substituting our value of ‘$\sigma$’ (the Prandlt number). Their formulation includes the role of horizontal turbulence rather than frictional drag. However, a value of $\sigma$ is not specified and is left to general interpretations. Spiegel and Zahn (1992) also have some estimates for the Sun’s current degree of penetration in the absence of drag, which will be discussed in the next section.
@@ 4.2 Time-dependent penetration
Figure 3: Time-dependent analytical and numerical solutions for the evolution of zonal velocity contours at early stages (till $t=10^{19}s$). The zonal wind forcing $u_{F}$ at the radiative-convective interface ($z=0$) is $10cm/s$. The boundary condition at the bottom is free shear i.e. $\partial u/\partial z=0$. The background colormap represents the magnitude of the zonal velocity, as obtained from the analytical solution, at different depths on the ordinate as time proceeds on the abscissa. Specific contours of the numerical Dedalus2 solution is shown using solid lines, while the dashed lines represent the analytical solution in addition to the background colormap. The zonal velocity is seen to decay on a timescale dictated by $k=5\times 10^{-19}$. This is followed by the slower secondary evolution in figure 4. Note: Maximum value of the zonal velocity at $z=0$ is NOT 10. This is a limitation of the analytical solution.
Figure 4: Figure 3 repeated for a longer time extending to $t=10^{21}s$ showing both the primary and secondary stages of penetration of the zonal wind. The contour and background colormap description is the same as Figure 3. The fast initial penetration, due to radiative diffusion, is followed by a slower phase over a timescale $\tau_{F}=3\times 10^{20}s$. The later occurs when the Rayleigh drag slows the radiative penetration down, finally bringing it to a halt. The zonal velocity penetrates through the entire depth (see contour corresponding to magnitude 1) during this stage owing to a weaker drag. Note: The analytical solution does well at later times.
@@
Previous time-dependent studies include Elliott (1997), which numerically solves a set of equations close to Spiegel and Zahn (1992) in spherical coordinates, although reporting results only till one solar age; Wood and Brummell (2012, 2018) perform full-blown three-dimensional numerical simulations of the radiative-convective interface including gravity waves, convective penetration and magnetic fields. They claim that the leading order dynamics are well captured by “balanced” models and may not require complex simulations. This sets the stage for the results obtained from our model, which, as discussed below, provide insights not discussed previously, using a significantly cheaper alternative and succinct explanation.
@@
Figures 3 and 4 compare the time-dependent analytical solutions (cf. expression (17)) with numerical solutions obtained using Dedalus (Burns et al., 2020). The temporal evolution of the zonal velocity is represented as contours on a spacetime plot at early and late stages, respectively. The value of $k=5\times 10^{-19}\hskip 2.84526ptcm^{2}/s$ and $\tau_{F}=10^{13}\hskip 2.84526ptyrs=3\times 10^{20}s$. The dashed and solid lines represent the analytical solution (also the background colormap) and Dedalus, respectively. The different intensities of the zonal wind clearly shows the penetration of the meridional circulation.
@@
Two distinct evolutionary phases are observed: one fast and the other slow. Neglecting the effect of rotation, stratification, and horizontal and vertical length scales, the aforementioned phases correspond to two competing effects determined by the magnitudes of $\tau_{R}$ and $\tau_{F}$. A strong radiative diffusion renders relatively fast penetration. The zonal velocity diffuses to a certain depth in time $\sim\tau_{R}$ while obeying the thermal wind balance. The Rayleigh drag, acting weakly, develops gradually and slows the penetration down until it comes to a complete halt on the timescale $\sim$ $\tau_{F}$. At steady-state radiative diffusion and Rayleigh drag balance each other out. This distinctive evolution is clear in the weak drag limit, $\tau_{R}<\tau_{F}$, as shown in figures 3 and 4. Thus, the time to steady-state is dictated by the longer drag time in this case. The internal rotation profile may therefore never reach steady-state if the drag is too weak, making the time-dependent solution directly relevant. However, in the limit of stronger drag i.e. $\tau_{R}\gtrsim\tau_{F}$, there is simply one evolutionary phase with hardly any penetration.
@@
Figure 5 (right panel) shows the time evolution of zonal velocity in the depths of the Sun, which, under all simplifying assumptions of the current work, must reach steady-state in $3\times 10^{20}s$ ($10^{13}\hskip 2.84526ptyrs$) or approximately 2000 times its current age. Hence, the Sun should run out of hydrogen fuel well before reaching rotational steady-state unless stronger drag is present. The left panel in Figure 5 compares the current and steady-state zonal velocity profiles using our formalism. The steady-state zonal/meridional flow penetrates all the way to the bottom of the radiative zone as also pointed out in Figure 2. In the current Sun, the tachocline extends through 0.75 pressure scale heights to a depth located at $0.625R_{\odot}$ from the center of the Sun.
@@
Figure 5: Left panel. Current ($4.6\times 10^{9}$ $yrs$, solid blue) and steady-state (solid red) zonal velocity profiles varying with depth in the radiative zone of the Sun. The dashed lines show the maximum depth of penetration of the zonal velocity. Right panel. Time evolution of the penetration of zonal wind in the Sun. The dash-dotted yellow vertical line represents the current age of the Sun.
@@
It is to be noted that the accuracy of the analytical solution is parameter dependent and breaks down at the boundaries because the $u=0$ initial condition throughout the domain cannot match the $u=u_{F}$ boundary condition. The discrepancy is more apparent at earlier times than late (see maximum colorbar values in both figures 3 and 4). The analytical solution also breaks down when $k$ and $1/\tau_{F}$ differ from one another by more than two orders of magnitude. Also, the linearized Rayleigh drag does not capture the scale dependence of an ordinary 2nd-order differential operator. Similarly, the radiative diffusion should ideally be replaced by a 4th-order hyperdiffusion term as in Spiegel and Zahn (1992). Despite these hefty caveats, the GFD-inspired thin-layer $f$-plane solution captures the leading order penetration phenomena with significant accuracy. It can be used to prototype differential rotation profiles in stellar interiors quickly, thereby allowing probes into the onset of magneto/hydrodynamic instabilities(waves) that potentially transport angular momentum in radiative envelopes of main-sequence stars. However, shortcomings of the analytical solution, variable parameters and hyperdiffusion call for detailed numerical solutions.
@@ 5 Conclusions
Our steady-state meridional streamfunction reaffirms the established understanding in stellar physics that the circulation decays over a length scale of $(N/2\Omega\times\sqrt{Pr})H$. Rotation and radiative diffusion are drivers of deep penetration, while a stable stratification and viscous drag act to oppose the effect. The differential rotation, which is a combination of the zonal wind and meridional circulation, propagates deep following a fast adiabatic adjustment to the thermal wind balance. We derive a time-dependent, one-dimensional, inhomogeneous diffusion equation which reasonably approximates the fourth-order hyperdiffusion process (Spiegel and Zahn, 1992) on a scale comparable to the pressure scale height. Closed-form analytical solutions allow rapid prototyping of differential rotation profiles at any time during the evolution of a main-sequence star.
Spiegel and Zahn (1992) have previously built a steady-state model for the penetration of differential rotation by invoking a strong turbulent (horizontal-only) viscosity to match helioseismic profiles. Using our time-dependent formalism, we also find that a sufficiently strong drag is required to truncate the penetration and reach steady state, otherwise, the internal rotation of main-sequence stars may keep evolving through their lifetime.
@@ Acknowledgements
KM is supported by the National Science and Engineering Research Council of Canada. This work has made extensive use of the following software packages: numpy, matplotlib, Dedalus.
We would furthermore like to acknowledge that this work was performed on land which for thousands of years has been the traditional land of the Huron-Wendat, the Seneca and the Mississaugas of the Credit. Today this meeting place is still the home to many indigenous people from across Turtle Island and we are grateful to have the opportunity to work on this land.
@@
The data underlying this article is available at: https://github.com/Textydeep/Meridional-Circulation-Dedalus2.git.
@@ Appendix A Hyperdiffusion
Here, $\kappa$ is the coefficient of radiative heat diffusion. Combining equations (11-13) and (18) we get
@@
Dimensional analysis shows that the regular 2nd-order diffusion evolves on a time scale of order $z^{2}\tau_{R}$. By contrast, the 4th-order hyperdiffusion operator evolves on a time scale of order $z^{4}\tau_{R}$. These two timescales match for $z=1$, i.e., on a scale equal to the pressure scale height $H$, which is also the scale we chose to relate $\tau_{R}$ and $\kappa$. In other words, we expect a regular diffusion treatment to approximate the actual hyperdiffusion process relatively well on scales of order H. We also note that the rather large scale height $H\sim 0.1R_{\odot}$ in a main sequence star like the Sun implies a modest dynamic range of only $z\sim 1\text{-}7$ over the full radiative zone.
@@
For all these reasons, we focus on a simpler second-order diffusion treatment in our main discussion to gain insight and obtain leading-order results. We recognize that numerical solutions to the exact hyperdiffusion equation 19 could be obtained relatively easily and would be necessary when the parameters in the hyperdiffusion equation vary with coordinates $z$. Nevertheless, we expect the main understanding gained from our second-order diffusion analysis with constant parameters to remain broadly valid and useful.

### v2

@@ 4 Applications
Let us consider a radiative zone that is initially under solid body rotation, $\varOmega$. At time $t=0$, a zonal wind, $u_{F}$, at the outer boundary ($z=0$) starts torquing this stably stratified layer. A fast, purely dynamical, adiabatic/geostrophic adjustment moves around fluid parcels creating local convergence and divergence at the surface (see figure 1c), thereby setting up a meridional circulation (refer §1), the depth of which depends on the ratio of rotation and stratification $\varOmega/N$, only. The circulation advects the background stratification to establish a meridional temperature gradient. Section 3 of Spiegel and Zahn (1992) gives a detailed description of the process.
Our solution aims to model the subsequent much slower evolution. The meridional temperature gradient developed in the adjusted layer (see Figure 1d panel 2) also impacts the vertical temperature profiles. Following a rapid thermal adjustment occurring on a short thermal timescale (Spiegel and Zahn, 1992), radiative diffusion propagates the horizontal temperature inhomogeneities further down; Figure 1d panel 3. As a result, the zonal wind and the meridional circulation spread deeper into the radiative envelope (Section 4 Spiegel and Zahn (1992)). The Rayleigh frictional drag (or viscosity) opposes this by acting on the vertical shear of the zonal wind. In steady-state, the meridional circulation penetrates to a depth until these two effects balance each other out. This slower process depends on the Prandtl number in addition to rotation and stratification, as discussed in §2 for the steady-state meridional streamfunction.
@@ 4.1 Steady-state zonal velocity
Here, $k$ accounts for the radiative timescale, rotation, and stratification via $\tau_{R}$, $f$, and $N$, respectively. Figure 2 shows vertically varying steady-state zonal wind profiles for different values of $k$ (black, constant $\tau_{F}=10^{13}\hskip 2.84526pt\rm{yrs}=3\times 10^{20}\rm{s}$) and $\tau_{F}$ (red, constant $k=5\times 10^{-19}\hskip 2.84526pt{\rm{cm}}^{2}/{\rm{s}}$) using expression (17). These are values adopted as representative of the Sun (Menou et al., 2004; Menou and Le Mer, 2006), the zonal velocity profile for which is labeled in the figure (black dashed, red solid lines). The radiative-convective interface is at $z=0$, and the bottom of the domain is at $z=\ ln(10^{3})=7$, considering the passive radiative zone to extend through pressures varying three orders in magnitude. The forcing zonal velocity, $u_{F}$, is $10\hskip 2.84526ptcm/s$ without loss of generality. The bottom boundary shears freely. Thus, there is no hard constraint on the velocity, allowing it to assume any value.
@@
From the same figure, the zonal wind is seen to penetrate all the way through the radiative zone of the Sun at steady-state. This is in reasonable agreement with the steady-state estimate made using equation (5.19) of Spiegel and Zahn (1992) by substituting our value of the Prandtl number. Their formulation includes the role of horizontal turbulence rather than frictional drag. However, a value for the Prandtl is not specified and is left to general interpretations. Spiegel and Zahn (1992) also have some estimates for the Sun’s current degree of penetration in the absence of drag, which will be discussed in the next section.
Figure 2: Analytical solution for steady-state zonal wind varying with the vertical coordinate indicating the depth of penetration of the meridional circulation. The top and bottom of the radiative zone are indicated by $z=0$ and $z=7$, respectively. Different values of $k$ (black) include the effects of rotation $f$, stratification $N$, meridional wavenumber $l$, scale height $H$ and the radiative relaxation timescale $\tau_{R}$ according to $k=1/(L_{D}^{2}l^{2}\tau_{R})$. See §3 for details. The effect of the viscous drag timescale $\tau_{F}$, on the depth of penetration is shown by red lines. The black dashed line overlapping the red solid line, on the left, depicts the steady-state zonal velocity profile in the solar radiative zone.
@@ 4.2 Time-dependent penetration
Previous time-dependent studies include Elliott (1997), which numerically solves a set of equations close to Spiegel and Zahn (1992) in spherical coordinates, although reporting results only till one solar age; Wood and Brummell (2012, 2018) perform full-blown three-dimensional numerical simulations of the radiative-convective interface including gravity waves, convective penetration and magnetic fields. They claim that the leading order dynamics are well captured by ‘balanced’ models and may not require complex simulations. This sets the stage for the results obtained from our model, which, as discussed below, provide a significantly cheaper alternative and succinct explanation.
@@
We use Dedalus2 (Burns et al., 2020) to numerically solve equation (16) and support the analytical solution. Dedalus is a pseudo-spectral code used to solve initial, boundary, or eigen-value problems with the flexibility of specifying the equations symbolically. A Chebyshev basis is used to discretize our 1D domain and Dirichlet and Neumann boundary conditions are applied at the top ($\bar{u}(0,t)=u_{F}$) and bottom ($u_{z}(L,t)=0$), respectively. Post discretization the system is evolved through the SBDF2 timestepper. Figure 3 compares the time-dependent analytical solutions (cf. expression (18)) with numerical solutions. The temporal evolution of the zonal velocity is represented as contours on a spacetime plot at early and late stages, respectively. The value of $k=5\times 10^{-19}\hskip 2.84526pt{\rm{cm}}^{2}/{\rm{s}}$ and $\tau_{F}=10^{13}\hskip 2.84526pt\rm{yrs}=3\times 10^{20}\rm{s}$. The dashed and solid lines represent the analytical solution (also the background colormap) and Dedalus, respectively. The different intensities of the zonal wind clearly shows the penetration of the meridional circulation.
Figure 3: Time-dependent analytical and numerical solutions for the evolution of zonal velocity contours at early stages (till $t=10^{19}s$). The zonal wind forcing $u_{F}$ at the radiative-convective interface ($z=0$) is $10cm/s$. The boundary condition at the bottom is free shear i.e. $\partial u/\partial z=0$. The background colormap represents the magnitude of the zonal velocity, as obtained from the analytical solution, at different depths on the ordinate as time proceeds on the abscissa. Specific contours of the numerical Dedalus2 solution is shown using solid lines, while the dashed lines represent the analytical solution in addition to the background colormap. The left and right panels show early and late stages of penetration, respectively. Note: Maximum value of the zonal velocity on the colorbar is NOT 10 despite the imposed BC, especially earlier (left panel). This is a limitation of the analytical solution.
@@
Two distinct evolutionary phases are observed: one fast and the other slow. If we assume rotation, stratification, and horizontal and vertical length scales to be constants, the aforementioned phases can be explained by two competing effects determined by the magnitudes of $\tau_{R}$ and $\tau_{F}$. A strong radiative diffusion renders relatively fast penetration. The zonal velocity diffuses to a certain depth in time $\sim\tau_{R}$ while obeying the thermal wind balance. The Rayleigh drag, acting weakly, develops gradually and slows the penetration down until it comes to a complete halt on the timescale $\sim$ $\tau_{F}$. At steady-state radiative diffusion and Rayleigh drag balance each other out. This distinctive evolution is clear in the weak drag limit, $\tau_{R}<\tau_{F}$, as shown in figure 3. Thus, the time to steady-state is dictated by the longer drag time in this case. The internal rotation profile may therefore never reach steady-state if the drag is too weak, making the time-dependent solution directly relevant. However, in the limit of stronger drag i.e. $\tau_{R}\gtrsim\tau_{F}$, there is simply one evolutionary phase with hardly any penetration (not shown here; can be reproduced with the Dedalus script shared at the end).
@@
Figure 4 (right panel) shows the time evolution of zonal velocity in the depths of the Sun, which, under all simplifying assumptions of the current work, must reach steady-state in $3\times 10^{20}\rm{s}$ ($10^{13}\hskip 2.84526pt\rm{yrs}$) or approximately 2000 times its current age. This is greater than both the Eddington-Sweet time and nuclear time for the Sun, which have values of $2.2\times 10^{11}\rm{yrs}$(Spiegel and Zahn, 1992) and $10^{10}\rm{yrs}$, respectively. Hence, the Sun should run out of hydrogen fuel well before reaching rotational steady-state unless stronger drag is present. The left panel in Figure 4 compares the current and steady-state zonal velocity profiles using our formalism. The steady-state zonal/meridional flow penetrates all the way to the bottom of the radiative zone as also pointed out in Figure 2. In the current Sun, the tachocline extends through 0.75 pressure scale heights to a depth located at $0.625\rm{R}_{\odot}$ from the center of the Sun.
@@
Figure 4: Left panel. Current ($4.6\times 10^{9}$ $\rm{yrs}$, solid blue) and steady-state (solid red) zonal velocity profiles varying with depth in the radiative zone of the Sun. Middle and right panels. Time evolution of the penetration of zonal wind in the Sun. The dash-dotted yellow vertical line in the middle panel corresponds to the current solar age, while the extreme right edge of the right panel corresponds to steady-state.
It is to be noted that the accuracy of the analytical solution is parameter dependent and breaks down at the boundaries because the $u=0$ initial condition throughout the domain cannot match the $u=u_{F}$ boundary condition. The discrepancy is more apparent at earlier times than late (see maximum colorbar values in figure 3). The analytical solution also breaks down when $k$ and $1/\tau_{F}$ differ from one another by more than two orders of magnitude. Also, the linearized Rayleigh drag does not capture the scale dependence of an ordinary 2nd-order differential operator. Similarly, the radiative diffusion should ideally be replaced by a 4th-order hyperdiffusion term as in Spiegel and Zahn (1992). Despite these hefty caveats, the GFD-inspired thin-layer $f$-plane solution captures the leading order penetration phenomena with significant accuracy. It can be used to prototype differential rotation profiles in stellar interiors quickly, thereby allowing probes into the onset of magneto/hydrodynamic instabilities(waves) that potentially transport angular momentum in radiative envelopes of main-sequence stars. However, shortcomings of the analytical solution, variable parameters and hyperdiffusion call for detailed numerical solutions.
@@ 5 Conclusions
Our steady-state meridional streamfunction reaffirms the established understanding in stellar physics that the circulation decays over a length scale proportional to $N/f\times\sqrt{Pr}$. Rotation and radiative diffusion are drivers of deep penetration, while a stable stratification and viscous drag act to oppose the effect. The differential rotation, which is a combination of the zonal wind and meridional circulation, propagates deep following a fast adiabatic adjustment to the thermal wind balance. We derive a time-dependent, one-dimensional, inhomogeneous diffusion equation which reasonably approximates the fourth-order hyperdiffusion process (Spiegel and Zahn, 1992) on a scale comparable to the pressure scale height. Closed-form analytical solutions allow rapid prototyping of differential rotation profiles at any time during the evolution of a main-sequence star.
Spiegel and Zahn (1992) have previously built a steady-state model for the penetration of differential rotation by invoking a strong turbulent (horizontal-only) viscosity to match helioseismic profiles. Using our time-dependent formalism, we also find that a sufficiently strong drag is required to truncate the penetration and reach steady-state, otherwise, the internal rotation of main-sequence stars may keep evolving through their lifetime. This could have potential consequences for chemical mixing and transport of angular momentum through meridional circulation and hydrodynamic/hydromagnetic instabilities in stellar interiors (Busse, 1981).
@@ Acknowledgements
DB is supported by the Department of Physics at University of Toronto. KM is supported by the National Science and Engineering Research Council of Canada. The authors would like to thank the anonymous reviewers for their reviews which significantly enhanced the paper.
@@
The data underlying this article is available at: https://github.com/Textydeep/Meridional-Circulation-Dedalus2.git
@@ Appendix A Hyperdiffusion
Here, $\kappa$ is the coefficient of radiative heat diffusion. Combining equations (12-14) and (19) we get
@@
Dimensional analysis shows that the regular 2nd-order diffusion evolves on a time scale of order $z^{2}\tau_{R}$. By contrast, the 4th-order hyperdiffusion operator evolves on a time scale of order $z^{4}\tau_{R}$. These two timescales match for $z=1$, i.e., on a scale equal to the pressure scale height $H$, which is also the scale we chose to relate $\tau_{R}$ and $\kappa$. In other words, we expect a regular diffusion treatment to approximate the actual hyperdiffusion process relatively well on scales of order H. We also note that the rather large scale height $H\sim 0.1\rm{R}_{\odot}$ in a main sequence star like the Sun implies a modest dynamic range of only $z\sim 1\text{-}7$ over the full radiative zone.
@@
For all these reasons, we focus on a simpler second-order diffusion treatment in our main discussion to gain insight and obtain leading-order results. We recognize that numerical solutions to the exact hyperdiffusion equation (20) could be obtained relatively easily and would be necessary when the parameters in the hyperdiffusion equation vary with coordinates $z$. Nevertheless, we expect the main understanding gained from our second-order diffusion analysis with constant parameters to remain broadly valid and useful.